I don't understand where they are getting the 5 bikes, 35 bikes and 12 bikes. I underlined the parts I'm confused about. If they were to ask you for net income or gross profit from this how would you calculate it? Thanks!First-In, First-Out First-in, first-out (FIFO sales occur, the costs of the earlie the costs from them first-out (FIFO) assum
Answer #1

I Have explained clearly If you have any doubts regardibg this answer please comment below

FIFO A Augs Begining Inventory = 10 + 891 910 Puchares = 154106 = 1590 closing 4 - 25 2500 sales sourits = 10x al = 910 Jo lo